OpenStudy (anonymous):

The Young's modulus and density of a certain kind of stainless steel are 180 GPa and 8 g cm 3 − respectively. (a) Estimate the speed of sound wave in the steel.

OpenStudy (anonymous):

"The value of v for waves traveling along the length of a bar or rod is defined by the Young's modulus and the density: \[v=\Bigg( \frac{Y}{\rho}\Bigg)^{1/2}\]" p.210 Vibrations and Waves, A.P. French. The derivation for it is on p. 172, from \[ \frac{\partial^2 \xi}{\partial x^2} = \frac{\rho}{Y} \frac{\partial^2 \xi}{\partial t^2} = \frac{1}{v^2}\frac{\partial ^2 \xi}{\partial t^2}\] \[v^2 = \frac{Y}{\rho}\] Where xi is longitudinal displacement from equilibrium - it's just the form of the wave equation applied to weird stress forces.

OpenStudy (anonymous):

thanks!!